OpenStudy (noahbred):

Please help If an object with a mass of 6 kg is pushed with a force of 42 N, at what rate will this object accelerate?

OpenStudy (agreene):

Remember that Newtons are defined \(kg\times m/s^{2}\) thus: \[42N = 6 kg \times a~~m/s^2\] solve for a.

OpenStudy (noahbred):

umm? 6^2